# CrEIMBO: Cross Ensemble Interactions in Multi-view Brain Observations

### v1

@@ Abstract
Modern recordings of neural activity provide diverse observations of neurons across brain areas, behavioral conditions, and subjects—thus presenting an exciting opportunity to reveal the fundamentals of brain-wide dynamics underlying cognitive function. Current analysis methods, however, often fail to fully harness the richness of such data as they either provide an uninterpretable representation (e.g., via “black box” deep networks) or over-simplify the model (e.g., assume stationary dynamics or analyze each session independently). Here, instead of regarding asynchronous recordings that lack alignment in neural identity or brain areas as a limitation, we exploit these diverse views of the same brain system to learn a unified model of brain dynamics. Specifically, we assume that brain observations stem from the joint activity of a set of functional neural ensembles (groups of co-active neurons) that are similar in functionality across recordings, and propose to discover the ensemble structures and their non-stationary dynamical interactions in a new model we term CrEIMBO (Cross-Ensemble Interactions in Multi-view Brain Observations). CrEIMBO identifies the composition of the per-session neural ensembles through graph-driven dictionary learning and models the ensemble dynamics as a latent sparse time-varying decomposition of global sub-circuits—thereby capturing non-stationary dynamics. CrEIMBO identifies multiple co-active sub-circuits while maintaining representation interpretability due to sharing sub-circuits across sessions. Furthermore, CrEIMBO distinguishes session-specific from global (session-invariant) computations by exploring when distinct sub-circuits are active. We demonstrate CrEIMBO’s ability to recover ground truth components in synthetic data and uncover meaningful brain dynamics, capturing cross-subject and inter- and intra-area variability, in high-density electrode recordings of humans performing a memory task.
Figure 1: Challenges and approach. Top right: Representation of the latent $K$ neural ensemble sub-circuits ($\{\bm{f}_{k}\}$) whose time-changing contributions to the full dynamics ($\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}$) governs the whole-brain activity. These subcircuits can either capture global task-related patterns (e.g. $\bm{f}_{1}$), capture general brain computations that vary over time and between sessions (e.g. $\bm{f}_{2}$), be active only under a specific session (e.g. $\bm{f}_{3}$), or capture both task related and general per-session computations at different period ($\bm{f}_{k}$, e.g., processing related to some visual cue presented at the beginning of the session only ). Top Left: Membership of neurons within functional neural ensembles. Bottom: Real-world data often provide distinct partial views (“sessions”) of the brain system that cannot be matched in terms of neural identity, quantity, or distribution between areas. This discrepancy may lead to biased conclusions, e.g., session 2 and session D lack representations from areas $a$ and $d$, respectively.
@@ 1 Introduction
Individual recording sessions provide a singular perspective of the brain dynamics, as the reinsertion of recording devices results in the capture of different subsets of neurons in each session. Thus fitting population-level models often reverts to per-session analysis (i.e., fitting each session independently), which is inherently constrained by the inability to incorporate the full set of recorded activity across all recorded brain areas across all sessions. Moreover, analyzing the data session-by-session, may hinder our ability to distinguish between computations that are session or subject specific from task related or global session invariant neural processes.
@@
Compounding the difficulty of merging different sessions’ data are the non-linear and non-stationary nature of neural activity, as well as the presence of noise and variability across trials—both of which require additional model complexity to account for. Moreover to extract scientifically meaningful insight, the models must be interpretable in that model parameters must be readily relatable to task variables or connections between recorded units. This gap between the opportunities offered by modern neural data and the limited capabilities of current methods necessitates new approaches to leverage the richness in modern brain data and discover the fundamental neural sub-circuits governing brain activity.
@@
One current approach used to combine multi-session data into a single model is the use of deep learning models. By training to predict all sessions’ data, deep learning models implicitly merge the datasets into the fit network weights. While these powerful models can combine information across sessions, their uninterpretable ‘black box’ nature and complexity typically hinders their ability to reveal the fundamental building blocks of neural computation. Other approaches, including tensor factorization and its variations (e.g., [3, 9]) may be limited in their ability to merge sessions of varying durations or identify the interactions between the identified neural components.
Here, we re-frame cross-session variability in neurons recorded and experimental settings not as a drawback, but as a valuable advantage. Rather than viewing the different sessions as capturing distinct neural populations, we perceive such data as providing multiple, complementary perspectives of one brain system. This approach enables a more thorough and holistic discovery of the underlying system designed to extract joint information from the entire dataset.
In particular, we hypothesize that the population dynamics at each session is driven by a common set of global sub-circuit interactions shared across sessions. To capture the non-linear and non-stationary aspects of the dynamics, we leverage recent advances in decomposed linear dynamical systems (dLDS), which enables us to model how the neural trajectories transverse a low-dimensional manifold via a time-varying combination of the sub-circuits. While we assume that the representation of the latent dynamics is shared across sessions, the measured neural ensembles are not. We thus model each session’s ensembles as independent neural representations, i.e., the membership of neurons to ensembles, that captures how the latent trajectories project into each sessions’ view of the brain.
@@
In this work we lay out this novel model, which we term Cross Ensemble Interactions in Multi-view Brain Observations (CrEIMBO), and demonstrate its ability to capture latent dynamics in multi-view neural data. Specifically our contributions include:
@@
  • We distinguish between intra- and inter-region interactions by identifying sparse, same-area neural ensembles, deploying a specialized structured prior over per-session projections from circuit space to neural space.
  • We accurately recover ground truth components in synthetic data and discover meaningful cross-regional brain interactions underlying human brain data utilizing a shared low-dimensional circuit space across recording sessions.
@@ 2 Related Work
Merging non-simultaneous neural recording sessions. Existing work for combining non-simultaneous neural population recordings, here termed “sessions”, include  [26] who introduced the term “stitching” for this problem and modeled neural activity with a shared latent dynamical systems model and [24] who proposed to stitch asynchronous recordings for inferring functional connectivity by assuming shared cross-session latent input and demonstrated the advantage in observing more neurons for inferring connectivity. However, none of these methods capture multiple ($>2$) interacting neural sub-circuits active simultaneously spanning populations across multiple areas, nor do they distinguish between within- and between-area interactions, nor do they address temporal non-stationarity. In [1, 2], Bishop et al. proposed integrating the neural activity structure into inference and expanding the proposed applications of stitching to study of communication between two neural populations of brain areas and proposed a theoretical framework of when stitching is possible. In [20], the authors proposed extracting low-dimensional dynamics from multiple sessions by learning temporal co-variations across neurons and time. Both of these methods, through, assume at least a small overlap in the identity of cross-session observed neurons. LFADS [11, 25] enables stitching while addressing the need for neural overlap by finding a shared non-linear dynamical model. However, the latent dynamics may be uninterpretable in terms of interactions among neural populations, and due to its foundation on a variational auto-encoder, it may be challenging to tune.
@@
Maintaining Interpretability in Non-Stationary Neural Dynamics Models. Real-world neural activity follows non-stationary nature due to changing environmental settings. Classical dynamical models often capture non-stationarity through various methods such as ’black box’ deep learning models (e.g. [22, 25, 29], which are powerful but are often not interpretable with respect to neural interactions; linear models, which are interpretable but not expressive enough to capture the complex nature of neural data; or switching piece-wise linear models  [15, 14, 19], which fail to enable the identification of co-active dynamics. dLDS (decomposed Linear Dynamical Systems) [16] (with its improved learning procedure in [18]), addresses this gap by generalizing the switching assumption to sparse time-changing decomposition of linear dynamical systems, thus enabling capturing co-active sub-circuits. However, while the current dLDS framework can be modified for co-processing multi-subject recordings that present variability in neuron identity and distribution, this modification lacks the structured regional prior over the loading matrix that is necessary to partition multi-area recordings such that between and within area interactions can be isolated and studied.
@@
Identification of Functional Neural Ensembles A potential way to find interpretable loading matrix that links low to high dimensional space is through the identification of functional neural ensembles—here refereed to as groups of neurons with similar functionalities—that determine the axes in the latent space. For instance, the mDLAG model used Automatic Relevance Determination (ARD) to promote population-wise sparsity patterns of each latent variable. In [17], the authors presented an approach to address sparsity in loading matrices while accounting for multi-trial and task-conditions variability. They proposed a graph-driven regularization technique that facilitates the grouping of co-active neurons into the same ensembles and pushing apart neurons that present different functionality, while allowing for structural adjustments in ensemble configurations across different task conditions. They further enable per-trial variability in temporal activity of the ensembles. However, their model does not model the dynamic interactions between ensembles or give a closed form prior for the evolution of the ensembles’ temporal traces. While other methods, e.g., clustering approaches like  [8] or Sparse Principal Component Analysis (SPCA) [30], enable the recognition of underlying sparse groups, they either do not allow a neuron to belong to multiple groups with varying degrees of membership or do not support controlled adaptations of structural membership across sessions.
@@
Multi-regional brain interactions. With the ability to record from multiple brain areas and the recognition of the importance of considering multi-area communication [21], recent works have delved into understanding the communication between or within brain areas. In [7, 6], the authors proposed models to identify time delays between cross-regional brain interactions, overcoming the challenge of distinguishing indirect, concurrent, and bi-directional interactions in two or more populations. However, their model focuses on analyzing sessions individually and is not intended to uncover the full underlying set of transition matrices, but rather to recognize meaningful time delays that can imply interactions. Other models addressed multi-regional interactions through a communication subspace [23] using dimensionality reduction, or by Generalized Linear Models (GLMs), with either Poisson [28] or Gaussian [28] statistics to identify functional coupling between areas. These approaches, though, are not tailored to capture non-simultaneous cross-sessions variability in neuron identity and subject processing. Other recent methods, including [10, 5, 13] discussed the potential usage of their model to multi-regional interactions, have proposed studying multi-regional interactions using switching dynamical systems approaches, where the model switches between different states over time to capture distinct periods. However, this approach cannot disentangle interactions among multiple sub-circuits active simultaneously, nor can it identify interactions at multiple temporal resolutions that occur concurrently and distinguish between these or multiple co-processed processes whose encoding occurs gradually rather than through sharp transitions.
3 Problem Formulation and Approach
Problem Formulation
@@ 3 Problem Formulation and Approach
Let $\{\bm{Y}^{d}\}_{d=1}^{D}$ be a set of estimated firing rates over $D$ asynchronous sessions, indexed by ${d=1\dots D}$. The neural recordings of each $d$ session, $\bm{Y}^{d}\in\mathbb{R}^{N^{d}\times T^{d}}$ captures $T^{d}$ time observations of the activity of $N^{d}$ neurons from up to $J$ distinct brain areas, thus offering a partial view of the brain system. Note that sessions may refer to the same or different subjects, can vary in duration or number of neurons, i.e. $N^{d}$ and $T^{d}$ need not be equal to $N^{d^{\prime}}$ and $T^{d^{\prime}}$ for different sessions $d,d^{\prime}$, and may differ in the sub-set of brain areas (out of $J$) they capture. Hence, these cross-session data matrices cannot be aligned into a single data array, e.g., via a tensor—hindering the direct application of existing analyses as tensor factorization.
We assume that these brain observations reflect the hidden activity $\bm{X}^{d}\in\mathbb{R}^{p\times T^{d}}$ of $p<<N^{d}$ functional neural groups that evolve and interact over time. Each of these groups encompasses a sparse set of same-area neurons, while each neuron can belong to more than one group with varying degrees of cross-group membership, encoded in the values of $\bm{X}^{d}$. The low-dimensional latent space in which the groups interact is shared across sessions and projected to each session’s observed neurons via an unknown per-session projection $g^{d}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{N^{d}}$ that captures the groups’ compositions in the visible neurons, such that $\bm{Y}^{d}=g^{d}(\bm{X}^{d})$.
@@
The interactions between these functional groups follow a per-session non-linear and non-statioary dynamics $\bm{F}_{t}^{d}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$, where $\bm{X}^{d}_{t}=\bm{F}^{d}_{t}(\bm{x}^{d}_{t-1})$. As in general non-stationary systems cannot be fit without constraints, we follow the dLDS model and assume that these dynamics can be described by a set $\{\bm{f}_{k}\}_{k=1}^{K}$ of $K$ global time- and session-invariant “sub-circuits”. Here each sub-circuit ${\bm{f}_{k}\in\mathbb{R}^{p\times p}}$ represents a basis interaction between these groups that is reused by the system at different time-points throughout its trajectory. Each of these latent sub-circuits may capture either “global” session-invariant brain interactions and/or session- and subject-specific interactions. Moreover, these sub-circuits may be active simultaneously or intermittently, yielding a model that can flexibly fit neural trajectories that differ between sessions with a single underlying mechanism.
The problem we address in this model is identifying the set of unknown multi-regional sub-circuits, including their neuronal composition and non-stationary activation levels, from asynchronously collected observations.
CrEIMBO (Cross Ensemble Interactions in Multi-view Brain Observations)
@@
A naive approach to identifying the unknown sub-circuits involves first projecting each session’s data into a low-dimensional space and subsequently fitting the dynamics model in the low-dimensional space. However, this two-step approach can be sub-optimal as the dimensionality reduction projections may not prioritize the same dimensions needed to preserve dynamic fidelity. An additional advantage of simultaneously identifying the ensembles and dynamics is that in the cross-session model it automatically aligns the cross-session components based on the shared dynamics prior. Such alignment is unlikely given the sensitivity of the subspace ordering in dimensionality reduction methods to each session’s distinct statistics.
@@
Let $\bm{A}^{d}$ be a sparse matrix where each column, $\bm{A}^{d}[:,j]$, encodes the neuronal composition of the $j$-th ensemble in the $d$-th session, such that non-zero values within these columns represent the membership of neurons in the ensemble and their magnitudes reflecting the degree of membership. Since we aim to capture multi-regional interactions, we design the ensemble matrix $\bm{a}:=\bm{A}^{d}$ as a block diagonal matrix where the $j^{th}$ block, $\bm{a}^{j}\in\mathbb{R}^{n_{j}\times p_{j}}$, contains the neural ensembles for the $j$-th area. We denote $p_{j}$ as the number of ensembles within that area (note that $p=\sum_{j=1}^{J}p_{j}$) and $n_{j}:=N_{j}^{d}$ as the number of neurons in that area (i.e., the total number of neurons observed in session $d$ is $N^{d}=\sum_{j=1}^{J}n_{j}$).
@@
Furthermore, let $\tau:=\mathrm{T}_{m}^{d}$ indicate the time duration (i.e., the number of time-points) of the $m^{th}$ trial, for $m=1\dots M_{d}$, in session $d$. In each trial, the temporal activity for the $p$ ensembles are denoted as $\bm{x}:=\bm{X}_{m}^{d}\in\mathbb{R}^{p\times\tau}$, leading to a model where the observations $\bm{y}:=\bm{Y}_{m}^{d}$ in trial $m$ are assumed to arise from the joint activity of all these ensembles, up do a normally-distributed error $\epsilon\sim\mathcal{N}(0,\sigma)$, i.e., $\bm{y}=\bm{a}\bm{x}+\epsilon.$ The evolution of the ensembles’ activities ($\bm{x}$) reflect the latent interactions between ensembles, which we model with a non-stationaly linear dynamical system $\bm{x}_{t}=\bm{F}_{t}\bm{x}_{t-1}$ for $t=1\ldots\tau$. The time-changing transition matrix $\bm{F}_{t}\in\mathbb{R}^{p\times p}$ captures both non-linear and non-stationary brain activity. A key desired property of our model is to capture the overlapping activity of multiple sub-circuits of interacting ensembles. We thus follow the dLDS concept and model the interactions ($\bm{F}_{t}$) through a time-varying sparse decomposition ${\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}}$, where $\{\bm{f}_{k}\}_{k=1}^{K}$ represents the $K$ global ensemble interactions, and $\bm{c}_{kt}$ are the time-varying coefficients that represent the modulation of each interaction pattern over time.

### v2

# CREIMBO: Cross-Regional Ensemble Interactions in Multi-view Brain Observations

@@ Abstract
Modern recordings of neural activity provide diverse observations of neurons across brain areas, behavioral conditions, and subjects; presenting an exciting opportunity to reveal the fundamentals of brain-wide dynamics. Current analysis methods, however, often fail to fully harness the richness of such data, as they provide either uninterpretable representations (e.g., via deep networks) or oversimplify models (e.g., by assuming stationary dynamics or analyzing each session independently). Here, instead of regarding asynchronous neural recordings that lack alignment in neural identity or brain areas as a limitation, we leverage these diverse views into the brain to learn a unified model of neural dynamics. Specifically, we assume that brain activity is driven by multiple hidden global sub-circuits. These sub-circuits represent global basis interactions between neural ensembles—functional groups of neurons—such that the time-varying decomposition of these sub-circuits defines how the ensembles’ interactions evolve over time non-stationarily and non-linearly. We discover the neural ensembles underlying non-simultaneous observations, along with their non-stationary evolving interactions, with our new model, CREIMBO (Cross-Regional Ensemble Interactions in Multi-view Brain Observations). CREIMBO identifies the hidden composition of per-session neural ensembles through novel graph-driven dictionary learning and models the ensemble dynamics on a low-dimensional manifold spanned by a sparse time-varying composition of the global sub-circuits. Thus, CREIMBO disentangles overlapping temporal neural processes while preserving interpretability due to the use of a shared underlying sub-circuit basis. Moreover, CREIMBO distinguishes session-specific computations from global (session-invariant) ones by identifying session covariates and variations in sub-circuit activations. We demonstrate CREIMBO’s ability to recover true components in synthetic data, and uncover meaningful brain dynamics in human high-density electrode recordings, including cross-subject neural mechanisms as well as inter- vs. intra-region dynamical motifs. Furthermore, using mouse whole-brain recordings, we show CREIMBO’s ability to discover dynamical interactions that capture task and behavioral variables and meaningfully align with the biological importance of the brain areas they represent.
Figure 1: CREIMBO’s Illustration. A: Real-world multi-regional neural datasets consist of multiple ($D$) non-simultaneous recording sessions ($\{{\bm{Y}^{d}}\}^{D}_{d=1}$) that cannot be matched in terms of individual neurons’ identities, quantity, or cross-region distribution, resulting in neurons appearing only in certain sessions (4-pointed stars) or not at all (gray circles). This variability hinder our ability to draw unified conclusions of whole-brain neural computation, while analyzing sessions individually could lead to session-specific biased results. This challenge is further complicated by variability in session (or trial) duration, the non-stationarity of brain dynamics (even within a session), and the presence of multiple trials within each session. B $\rightarrow$ A: Instead of considering these asynchronous sessions as a challenge, we frame them as an opportunity to obtain distinct, potentially complementary “views” onto the shared underlying brain system, thereby facilitating the learning of a unified brain dynamics model. We assume that brain computations are mediated by multiple neural ensembles—$\bigstar$ groups of same-region neurons with shared functionalities—whose interactions yield meaningful neural representations. However, the specific neurons and their membership degrees within each ensemble are unknown, and a neuron may belong to multiple ensembles with varying membership levels based on its diverse functionalities. C: CREIMBO leverages partial brain views to co-learn ensemble compositions per session ($\{\bm{A}^{d}\}_{d=1}^{D}$), their temporal activity ($\{\bm{x}^{d}\}_{d=1}^{D}$), and cross-regional interactions $\bm{x}^{d}_{t+1}=\bm{F}_{t}^{d}\bm{x}^{d}_{t}$. It posits that these non-stationary interactions stem from a session-shared dictionary prior of up to $K$ global interactions $\{\bm{f}_{k}\}_{k=1}^{K}$, whose sparse time-changing decomposition $\bm{F}^{d}_{t}=\sum_{k=1}^{K}\bm{f}_{k}\bm{c}_{kt}^{d}$ shapes the overall ensemble dynamics at each time point $t$ and session $d$. The interactions’ coefficients can help distinguish session-specific interactions from session-invariant interactions in different time periods (pink right, solid vs. dotted in subplot C right). With the session-shared ensemble-interactions dictionary and certain assumptions (Sec. Assumptions), the model auto-sorts ensembles across sessions by functionality, aligning the $j$-th ensemble of session $d$ with the $j$-th ensemble of session $d^{\prime}$, despite differences in observed neurons.
@@ 1 Introduction
Individual recording sessions provide a singular perspective of the brain dynamics, as the reinsertion of recording devices results in the capture of different subsets of neurons in each session. Thus, fitting population-level models often reverts to per-session analysis (i.e., fitting each session independently), which is inherently constrained by the inability to incorporate the full set of recorded activity across all recorded brain areas across all sessions. Moreover, analyzing the data session-by-session, may hinder our ability to distinguish between computations that are session or subject specific from task related or global session-invariant neural processes. Compounding the difficulty of merging different sessions’ data are the non-linear and non-stationary nature of neural activity, as well as the presence of noise and variability across trials—both of which require additional model complexity to account for. Moreover, to extract scientifically meaningful insights, the models must be interpretable, with model parameters directly related to task variables or connections between recorded units. This gap between the opportunities offered by modern neural data and the limited capabilities of current methods necessitates new approaches to leverage the richness in modern brain data and discover the fundamental neural sub-circuits governing brain activity.
@@
One current approach used to combine multi-session data into unified model is deep learning. By training to predict all sessions’ data, deep networks implicitly merge the datasets into the fit network weights. While these powerful models can combine information across sessions, their uninterpretable nature and complexity typically hinders their ability to reveal the fundamental building blocks of neural computation. Other approaches, including tensor factorization and its variations (e.g., Chen et al. (2015); Harshman (1970)) may be limited in their ability to merge sessions of varying durations or model the interactions between the identified neural components.
Here, we re-frame cross-session variability in neurons recorded and experimental settings, not as a drawback, but as a valuable advantage, viewing this data as providing multiple, complementary “views” into a single, shared brain system (Fig. 1B $\rightarrow$ A). This approach enables a more thorough and holistic discovery of the underlying system designed to extract joint information from the entire dataset. In particular, we hypothesize that the population dynamics at each session is driven by a common set of global time-invariant sub-circuits shared across sessions. To effectively capture the non-linear and non-stationary dynamics of the brain, we enable these sub-circuits to adjust their activity levels over time and across sessions. This approach allows us to identify evolving neural patterns through the collective activity of the sub-circuits, modeled via their time-varying decomposition. We assume that the latent dynamics governed by functional neural groups are shared across sessions, even though the specific individual neurons observed may vary. To address this, we propose a per-session transformation from the joint low-dimensional circuit space to each session’s observation space, ensuring cross-session alignment of the neural groups’ functionalities.
@@
In this work, we lay out this new model, which we term Cross-Regional Ensemble Interactions in Multi-view Brain Observations (CREIMBO), and demonstrate its ability to capture latent dynamics in multi-view neural data. Specifically, our contributions include:
@@
  • We distinguish between intra- and inter-region interactions by identifying sparse, same-area neural ensembles, deploying a specialized structured prior over per-session projections from the circuit latent space to the observed neural space.
  • We accurately recover ground truth components in synthetic data and discover meaningful cross-regional brain interactions underlying human and mouse brain data.
@@ 2 Related Work
Merging non-simultaneous neural recording sessions. Existing work for combining non-simultaneous neuronal population recordings, here termed “sessions”, include  Turaga et al. (2013) who introduced the term “stitching” for this problem and modeled neural activity with a shared latent dynamical systems model and Soudry et al. (2013) who proposed to stitch asynchronous recordings for inferring functional connectivity by assuming shared cross-session latent input and demonstrated the advantage in observing more neurons for inferring connectivity. However, these methods are not designed to capture more than two simultaneously interacting neural sub-circuits across multiple populations, nor do they distinguish between within- and between-area interactions or specifically address temporal non-stationarity. Bishop (2015) and Bishop & Yu (2014) proposed integrating the neural activity structure into inference and expanding stitching to study the communication between two neuronal populations from potentially different brain areas. In Nonnenmacher et al. (2017), the authors proposed extracting low-dimensional dynamics from multiple sessions by learning temporal co-variations across neurons and time. Both of these methods, through, assume at least a small overlap in the identity of cross-session observed neurons. LFADS (Keshtkaran et al., 2021; Pandarinath et al., 2018) enables stitching while addressing the need for limited neural overlap by finding a shared non-linear latent dynamical model. However, its latent dynamics may be difficult to interpret regarding the specific neural sub-circuits that compose it, and due to its foundation on a variational auto-encoder, it may be challenging to tune.
@@
Maintaining Interpretability in Non-Stationary Neural Dynamics Models. Real-world neural activity follows non-stationary nature due to changing environmental settings. Classical dynamical models often simplified neural dynamics to linear (e.g., jPCA Churchland et al. (2012)) or non-linear but stationary (e.g., Gallego et al. (2020)). Models that capture non-stationarity often derive from “black box” deep learning models (e.g., Schneider et al. (2023); Pandarinath et al. (2018); Zhu et al. (2022), which are powerful but often present limited interpretability with respect to neural interactions. Other models are built on switching piece-wise linear models  Linderman et al. (2016; 2017); Murphy (1998), which do not enable the identification of multiple co-active dynamic processes. dLDS (decomposed Linear Dynamical Systems,  Mudrik et al. (2024a)) and its multi-step extension (Mudrik et al., 2024c)), addresses this gap by generalizing the switching assumption to sparse time-changing decomposition of linear dynamical systems, thus enabling capturing co-active sub-circuits. However, while dLDS can be extended to co-process multi-session recordings with cross-session variability in neuron identity, this modification may not yield interpretable latent-to-observed projections and lacks the structured multi-regional prior over the emission matrix necessary to partition multi-area recordings and study between- vs. within-area interactions.
@@
Identification of Functional Neural Ensembles. A potential way to find interpretable loading matrix that links low to high dimensional space is through the identification of functional neural ensembles—here referred to as groups of neurons with similar functionalities—that determine the axes in the latent space. For instance, the mDLAG model used Automatic Relevance Determination (ARD) to promote population-wise sparsity patterns of each latent variable. In Mudrik et al. (2024b), the authors presented an approach to address sparsity in loading matrices while accounting for multi-trial and task-conditions variability. They proposed a graph-driven regularization technique that facilitates the grouping of co-active neurons into the same ensembles and pushing apart neurons that present different functionality, while allowing for structural adjustments in ensemble configurations across different task conditions. They further enable per-trial variability in temporal activity of the ensembles. However, their model does not model the dynamic interactions between ensembles or give a closed form prior for the evolution of the ensembles’ temporal traces. While other methods, e.g., clustering approaches like  (Grossberger et al., 2018) or Sparse Principal Component Analysis (SPCA) Zou et al. (2006), enable the recognition of underlying sparse groups, they either do not allow a neuron to belong to multiple groups with varying degrees of membership or do not support controlled adaptations of structural membership across conditions.
@@
Multi-regional brain interactions. With the ability to record from multiple brain areas and the recognition of the importance of considering multi-area communication (Pesaran et al., 2021), recent works have delved into understanding the communication between or within brain areas. In Gokcen et al. (2022; 2024), the authors proposed models to identify time delays between cross-regional brain interactions, overcoming the challenge of distinguishing indirect, concurrent, and bi-directional interactions in two or more populations. However, their model focuses on analyzing sessions individually and is not intended to uncover the full underlying set of transition matrices, but rather to recognize meaningful time delays that can imply interactions. Other models addressed multi-regional interactions through a communication subspace Semedo et al. (2019) using dimensionality reduction, or by Generalized Linear Models (GLMs), with either Poisson (Yates et al., 2017) or Gaussian (Yates et al., 2017) statistics to identify functional coupling between areas. These approaches, though, are not tailored to capture non-simultaneous cross-sessions variability in neuron identity and subject processing. Other recent methods (Karniol-Tambour et al. (2024); Glaser et al. (2020); Li et al. (2024)) discussed the potential usage of their model to multi-regional interactions, have proposed studying multi-regional interactions using switching dynamical systems approaches, where the model switches between different states over time to capture distinct periods. However, this approach cannot disentangle interactions among multiple sub-circuits active simultaneously, nor can it identify interactions at multiple temporal resolutions that occur concurrently and distinguish between these or multiple co-processed processes whose encoding occurs gradually rather than via sharp transitions.
3 Problem Introduction and Approach
@@
Problem Intuition: Let $\{\bm{Y}^{d}\}_{d=1}^{D}$ be a set of estimated neural firing rates (e.g. from neuropixels data) over $D$ asynchronous sessions, indexed by ${d=1\dots D}$. The neural recordings of each $d$ session, $\bm{Y}^{d}\in\mathbb{R}^{N^{d}\times T^{d}}$, capture $T^{d}$ time observations of the activity of $N^{d}$ neurons from up to $J$ distinct brain areas, thus offering a partial view of the brain system. Note that sessions may refer to the same or different subjects, can vary in duration or number of neurons, i.e. $N^{d}$ and $T^{d}$ need not be equal to $N^{d^{\prime}}$ and $T^{d^{\prime}}$ for different sessions $d,d^{\prime}$, and can differ in the subset of brain areas (out of $J$) they capture. Hence, these cross-session data matrices cannot be aligned into a single data array, e.g., via a tensor—hindering the direct application of existing analyses as tensor factorization.
@@
We assume that these brain observations reflect the hidden activity $\bm{X}^{d}\in\mathbb{R}^{p\times T^{d}}$ of $p<<N^{d}$ functional neural groups that evolve and interact over time. Each of these groups encompasses a sparse set of same-area neurons, while each neuron can belong to more than one group with varying degrees of cross-group membership, encoded in the values of $\bm{X}^{d}$. The low-dimensional latent space in which the groups interact is shared across sessions and projected to each session’s observed neurons via an unknown per-session projection $g^{d}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{N^{d}}$ that captures the groups’ compositions in the visible neurons, such that $\bm{Y}^{d}=g^{d}(\bm{X}^{d})$. The interactions between these functional groups follow a per-session non-linear and non-statioary dynamics $\bm{F}_{t}^{d}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$, where $\bm{X}^{d}_{t}=\bm{F}^{d}_{t}(\bm{x}^{d}_{t-1})$. As in general non-stationary systems cannot be fit without constraints, we follow the dLDS model and assume that these dynamics can be described by a set $\{\bm{f}_{k}\}_{k=1}^{K}$ of $K$ global time- and session-invariant “sub-circuits”. Here each sub-circuit ${\bm{f}_{k}\in\mathbb{R}^{p\times p}}$ represents a basis interaction between these groups that is reused by the system at different time-points throughout its trajectory. Each of these latent sub-circuits may capture either “global” session-invariant brain interactions and/or session- and subject-specific interactions. Moreover, these sub-circuits may be active simultaneously or intermittently, yielding a model that can flexibly fit neural trajectories that differ between sessions with a single underlying mechanism. The problem CREIMBO addresses is thus identifying the set of $K$ latent multi-regional sub-circuits $\{\bm{f}_{k}\}_{k=1}^{K}$, capturing interactions between unknown functional sparse neural ensembles $\bm{A}$, along with their non-stationary activation levels $\{\{\bm{c_{kt}^{d}}\}_{k=1}^{K}\}_{d=1}^{D}$, by leveraging the joint information from asynchronously collected observations $\{\bm{Y}^{d}\}_{d=1}^{D}$.
@@
CREIMBO: A naive approach to identifying the unknown sub-circuits involves first projecting each session’s data into a low-dimensional space and subsequently fitting the dynamics model in the low-dimensional space. However, this two-step approach can be sub-optimal as the dimensionality reduction projections may not prioritize the same dimensions needed to preserve dynamic fidelity. An additional advantage of simultaneously identifying the ensembles and dynamics is that in the cross-session model it automatically aligns the cross-session components based on the shared dynamics prior. Such alignment is unlikely given the sensitivity of the subspace ordering in dimensionality reduction methods to each session’s distinct statistics.
@@
Let $\bm{A}^{d}$ be a sparse matrix where each column, $\bm{A}^{d}[:,j]$, encodes the neuronal composition of the $j$-th ensemble in the $d$-th session, such that non-zero values within these columns represent the membership of neurons in the ensemble and their magnitudes reflecting the degree of membership. Since we aim to capture multi-regional interactions, we design the ensemble matrix $\bm{a}:=\bm{A}^{d}$ as a block diagonal matrix where the $j$-th block, $\bm{a}^{j}\in\mathbb{R}^{n_{j}\times p_{j}}$, contains the neural ensembles for the $j$-th area. We denote $p_{j}$ as the number of ensembles within that area (note that $p=\sum_{j=1}^{J}p_{j}$) and $n_{j}:=N_{j}^{d}$ as the number of neurons in that area (i.e., the total number of neurons observed in session $d$ is $N^{d}=\sum_{j=1}^{J}n_{j}$). Furthermore, let $\tau:=\mathrm{T}_{m}^{d}$ indicate the number of time-points recorded in each $m$-th trial of session $d$ ($m=1\dots M_{d}$), such that the observations in that trial are captured by $\bm{Y}_{m}^{d}\in\mathbb{R}^{N^{d}\times\mathrm{T}_{m}^{d}}$. In each trial $m$, the temporal activity for the $p$ ensembles are denoted as $\bm{x}:=\bm{X}_{m}^{d}\in\mathbb{R}^{p\times\tau}$, leading to a model where the observations $\bm{y}:=\bm{Y}_{m}^{d}$ in trial $m$ are assumed to arise from the joint activity of all these ensembles, up do a normally-distributed error $\epsilon\sim\mathcal{N}(0,\sigma)$, i.e., $\bm{y}=\bm{a}\bm{x}+\epsilon.$ Please refer to Section G for Poisson statistics inference in low spiking rate regimes. The evolution of the ensembles’ activities ($\bm{x}$) reflect the latent interactions between ensembles, which we model with a non-stationary linear dynamical system $\bm{x}_{t}=\bm{F}_{t}\bm{x}_{t-1}$ for $t=1\ldots\tau$. The time-changing transition matrix $\bm{F}_{t}\in\mathbb{R}^{p\times p}$ captures both non-linear and non-stationary brain activity. A key desired property of our model is to capture the overlapping activity of multiple sub-circuits of interacting ensembles. We thus follow dLDS and model the interactions ($\bm{F}_{t}$) through a time-varying sparse decomposition ${\bm{F}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}}$, where $\{\bm{f}_{k}\}_{k=1}^{K}$ are the $K$ global ensemble interactions, and $\bm{c}_{kt}$ are their time-varying coefficients that capture the modulation of each interaction.